SPEED-DEPENDENT EFFECTS ON THE LINE SHAPES OF ARGON-BROADENED HF

\underline{A. S. PINE}, Alpine Technologies, 14401 Poplar Hill Road, Germantown, MD 20874.

Asymmetric line shapes have been observed for infrared transitions of Ar-broadened HF ^{a} and were fit to Dicke-narrowed profiles with an empirical partial correlation between velocity- and state-changing collisions. That analysis yielded measured broadening and shifting coefficients in excellent agreement with the thermally-averaged values calculated from exact quantum close-coupled scattering cross-section based on a realistic van der Waals potential ^{b}. However, some anomalies in the sense of the asymmetries and the magnitude and J dependence of the velocity-changing rates remained, which subsequently were addressed classically by Joubert et al. ^{c} and quantum mechanically by Demeio et al. ^{d}. Here, we reexamine those line shapes with the correlated speed-dependent strong-collision Dicke-narrowed profiles described above using the reliable energy-dependent cross-sections of Green and Hutson. For HF/Ar below one atmosphere, Dicke narrowing dominates the speed dependence.

aA. Pine, J. Chem. Phys. 101, 3444 (1994).

bS. Green and J. Hutson, J. Chem. Phys. 100, 891 (1994).

cP. Joubert, J. Bonamy, D. Robert, J. Domenech and D. Bermejo, JQSRT (to be published).

dL. Demeio, S. Green and L. Monchick, J. Chem. Phys. 102, 9160 (1995).